OpenStudy (anonymous):

Use the distributive property to simplify the expression below: one over two(g - 8)

OpenStudy (anonymous):

1 over 2 it is 1/2 ?

OpenStudy (anonymous):

yes,ok...

OpenStudy (anonymous):

1/2g-4

OpenStudy (anonymous):

you destribute the 1/2 to the terms inside the parenthesis. So g times 1/2 and -8 times 1/2. g*1/2 = 1/2g -8 * 1/2 (same as dividing by 2) = -4 1/2g - 4

OpenStudy (anonymous):

Vacuole's answer is right.